# New Paradigm of Adversarial Training: Breaking Inherent Trade-Off between Accuracy and Robustness via Dummy Classes

### v1

@@ Abstract
Adversarial Training (AT) is recognized as one of the most effective methods to enhance the robustness of Deep Neural Networks (DNNs). However, existing AT methods suffer from an inherent trade-off between adversarial robustness and clean accuracy, which seriously hinders their real-world deployment. Previous works have studied this trade-off within the current AT paradigm, exploring various factors such as perturbation intensity, label noise and class margin. Despite these efforts, current AT methods still typically experience a reduction in clean accuracy by over 10% to date, without significant improvements in robustness compared with simple baselines like PGD-AT. This inherent trade-off raises a question: whether the current AT paradigm, which assumes to learn the corresponding benign and adversarial samples as the same class, inappropriately combines clean and robust objectives that may be essentially inconsistent. In this work, we surprisingly reveal that up to 40% of CIFAR-10 adversarial samples always fail to satisfy such an assumption across various AT methods and robust models, explicitly indicating the improvement room for the current AT paradigm. Accordingly, to relax the tension between clean and robust learning derived from this overstrict assumption, we propose a new AT paradigm by introducing an additional dummy class for each original class, aiming to accommodate the hard adversarial samples with shifted distribution after perturbation. The robustness w.r.t. these adversarial samples can be achieved by runtime recovery from the predicted dummy classes to their corresponding original ones, eliminating the compromise with clean learning. Building on this new paradigm, we propose a novel plug-and-play AT technology named DUmmy Classes-based Adversarial Training (DUCAT). Extensive experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet demonstrate that the DUCAT concurrently improves clean accuracy and adversarial robustness compared with state-of-the-art benchmarks, effectively breaking the existing inherent trade-off. The code is available at https://github.com/FlaAI/DUCAT.
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated remarkable performance in various real-world applications, but they remain vulnerable to adversarial attacks (Biggio et al., 2013; Szegedy et al., 2014). Specifically, malicious input perturbations that are often imperceptible to humans can cause significant changes in the output of DNNs (Goodfellow et al., 2015; Huang et al., 2020), raising serious security concerns within both the public and research communities (Wang et al., 2020b). In response, several defense mechanisms have been proposed to enhance the adversarial robustness of DNNs, such as defense distillation (Papernot et al., 2017), feature squeezing (Xu et al., 2018), randomization (Xie et al., 2018), and input denoising (Guo et al., 2018; Liao et al., 2018). However, most of these techniques have proven subsequently to rely on obfuscated gradients (Athalye et al., 2018) and be ineffective against more advanced adaptive attacks (Tramer et al., 2020).
@@
Currently, Adversarial Training (AT) (Goodfellow et al., 2015; Madry et al., 2018) is demonstrated as one of the most effective approaches to train inherently robust DNNs (Athalye et al., 2018; Dong et al., 2020). Different from clean training, where Empirical Risk Minimization (ERM) serves as the fundamental paradigm, AT directly utilizes adversarially augmented samples to yield robust models. Research indicates that this paradigm is equivalent to optimizing an upper bound of natural risk on the original data, and can thereby serve as a principle against adversarial attacks (Tao et al., 2021). To the best of our knowledge, despite considerable advancements in the specific mechanism of existing AT methods, most of them remain adhering to this principal paradigm. The solid AT benchmarks such as PGD-AT (Madry et al., 2018), TRADES (Zhang et al., 2019) and MART (Wang et al., 2020b) are representative examples.
However, recent studies indicate that existing AT methods suffer from an inherent trade-off between adversarial robustness and clean accuracy (Tsipras et al., 2019; Zhang et al., 2019; Raghunathan et al., 2019; 2020; Wang et al., 2020a; Bai et al., 2021). This means improving robustness via these AT methods is at the cost of reducing model accuracy compared with the standard training, which can seriously hurt the experience of benign users and greatly reduce the use of AT among real-world DNN application providers. A widely recognized cause of this problem is that AT equally requires predictive consistency within the $\epsilon$-ball of each sample, which can complicate decision boundaries, particularly for those samples near class margins, ultimately degrading clean generalization (Dong et al., 2022; Rade & Moosavi-Dezfooli, 2022; Cheng et al., 2022; Yang & Xu, 2022). To learn robustness w.r.t. such samples more appropriately, FAT (Zhang et al., 2020) and HAT (Rade & Moosavi-Dezfooli, 2022) utilized adaptive perturbations in AT to smooth decision boundaries; Consistency-AT (Tack et al., 2022) encouraged the similarity of predictive distributions between adversarial samples derived from different augmentations of the same instance, thereby enhancing learnable patterns and reducing label noise; SOVR (Kanai et al., 2023) explicitly increased logits margin of important samples by switching from cross-entropy to a new one-vs-the-rest loss.
Despite several previous efforts, there has been limited substantial progress in addressing this trade-off problem. Most impressive advancements in recent years come from those methods introducing extra data (Alayrac et al., 2019; Carmon et al., 2019; Najafi et al., 2019; Zhai et al., 2019; Li et al., 2022) or utilizing generative models (Wang et al., 2023b). However, they not only demand more resources and computational costs but also violate the conventional fairness assumption of AT (i.e., no additional data should be incorporated (Pang et al., 2021)). Other advanced AT techniques apart from these still typically experience a drop in clean accuracy exceeding 10% to date, and the state-of-the-art (SOTA) robustness has just improved slowly (i.e., about 1% per year on average since 2018) (Wei et al., 2023; Dong et al., 2023; Li & Spratling, 2023; Jia et al., 2024). This inherent trade-off raises a question about the current AT paradigm on uniformly learning accuracy and robustness:
Figure 2: High overlap between adversarial samples escaping from the four AT benchmarks. This implies that such failures are more likely caused by an inappropriate learning objective from the overstrict assumption of the current AT paradigm, rather than specific ineffective AT methods.
Is the current AT paradigm, which compels DNNs to classify corresponding benign and adversarial samples into the same class, really appropriate and necessary for achieving adversarial robustness?
@@
In this work, as straightforward evidence supporting our deduction, we reveal that certain samples that always fail to meet the above objective of the current AT paradigm generally exist across various AT methods and different robust models. This observation implies that the conventional assumption of categorizing adversarial samples to the same class as the corresponding benign ones may be overstrict. As such, with the learning of benign and adversarial samples as two essentially inconsistent targets, the attempt to unify them in the current AT paradigm can be improper, which is likely to be blamed for the existing trade-off between clean accuracy and adversarial robustness. In response, we propose a new AT paradigm introducing additional dummy classes for certain adversarial samples that differ in distribution from the original ones to relax the current overstrict assumption. Then accordingly, we propose a novel AT method called DUmmy Classes-based Adversarial Training (DUCAT), breaking the inherent trade-off between clean accuracy and adversarial robustness.
@@
Our core idea is to create dummy classes with the same number as the original ones and respectively attribute them as the primary targets for benign and adversarial samples during training. Importantly, we do not suggest a strict separation between corresponding benign and adversarial samples (i.e., to make the original classes completely benign and the dummy ones completely adversarial), because assuming corresponding benign and adversarial samples have completely different distributions is still excessive, which may cause memorization of hard-label training samples, resulting in overfitting to specific adversaries as demonstrated by our toy case in Section 2.2.3. Instead, we construct soft labels to explicitly bridge corresponding original and dummy classes as the suboptimal alternative target of each other, so that the separation also becomes learnable, to utilize the potential of DNNs further. During inference time, such one-to-one correspondences enable the detection and recovery of adversarial samples, thereby achieving robustness without significantly compromising clean learning. Specifically, if a test sample is classified into a dummy class, this serves as an indication of a potential adversarial attack, allowing us to recover its clean prediction through a projection back to the corresponding original class. Separated from the computation graph of DNN and implemented in a run-time-only manner, such a projection further degrades the ability of real-world adversaries.
The main contributions of this work can be summarized as follows:
  • For the first time, we empirically reveal that always-failed samples widely exist in conventional AT, explicitly suggesting the assumption of the current AT paradigm to learn benign and adversarial samples with the same labels is overstrict;
  • We propose a new AT paradigm introducing dummy classes to relax the current assumption, breaking the inherent trade-off between accuracy and robustness from it; and
  • A novel plug-and-play DUCAT method is proposed, concurrently improving the accuracy and robustness of four common AT benchmarks in large-scale experiments on CIFAR-10, CIFAR-100 and Tiny-ImageNet, significantly outperforming 14 SOTAs.
@@ 2 Involving Dummy Classes in Adversarial Training
2.1 Motivation: Always-Failed Samples Widely Exist in AT
The undesirable progress regarding the trade-off between clean accuracy and adversarial robustness in the past years makes us suspect that the assumption of the current AT paradigm to assign corresponding benign and adversarial samples to the same class may be inappropriate and unnecessary. As intuitive evidence of this deduction, by two proof-of-concept experiments, we demonstrate that up to 40% adversarial samples that always fail to meet such an assumption generally exist in conventional AT crossing various AT methods and different robust models. The experiments are conducted on CIFAR-10 (Krizhevsky & Hinton, 2009) and ResNet-18 (He et al., 2016), with, as mentioned in Section 1, the most popular AT benchmarks, PGD-AT, TRADES, MART, and a representative SOTA, Consitency-AT. The training details are the same as our main experiments in Section 3.1. For adversary, we adopt a typical and effective adversarial attack, PGD-10 (Madry et al., 2018).
@@ 2.1 Motivation: Always-Failed Samples Widely Exist in AT
We first demonstrate that there is a high overlap in both successful and unsuccessful cases of the four AT methods. Specifically, for every benign test sample, with a high probability (i.e., $>$ 80%), its corresponding adversarial sample can either be uniformly defended by the four methods or evade all of them. As illustrated in Figure 2, we report the numbers of samples that can respectively beat none, one, two, three and all experimental defenses for different classes. It is clear that the samples in which all the defenses are uniformly worked or failed make up the overwhelming majority. Besides, the confusion matrixes of the models protected by these four AT methods under the adversary are shown in Figure 8 in Appendix C, where similar failure patterns can be noted crossing all of them. Based on these observations, we state that always-failed adversarial samples widely exist no matter the specific choice of AT methods.
Figure 3: The robust models already enhanced by different AT methods are still highly likely to be uniformly beaten by the successful adversarial samples generated based on any one of these models. Such a deconstruction of adversarial transferability between robust models reveals the model-independent vulnerability of certain samples, which further supports our deduction for the current AT paradigm.
@@
By respectively adopting the four AT benchmarks to train a model from scratch and then test it by the same white-box adversary, above we have revealed that the always-failed cases are independent of specific AT methods. In this part, through black-box transfer attacks, we further demonstrate that such cases should not be attributed to the weakness of any specific robust models as well. To be specific, we first train four different robust models respectively by the benchmarks and select any one of them as the surrogate model, based on which we generate adversarial samples for the transfer attack. Then we divide the generated samples into two subsets based on whether successfully attack the surrogate model, and respectively use them to attack the other three models. Figure 3 shows the results with the four robust models serving as the surrogate one in order, where significant differences in transferability between the originally successful and unsuccessful subsets can be clearly observed. In other words, for every single adversarial sample generated based on any robust model, if it is originally effective/ineffective in attacking this surrogate model, then with a high probability, it would also work/fail on other robust models. This implies that always-failed cases generally exist regardless of the specific robust model as the target.
@@
Now that the wide existence of always-failed samples is confirmed crossing different AT methods and robust models, compared with simply attributing this to specific technology or implementation details, it seems more reasonable to rethink if it is the paradigm to be blamed in essence. Based on the analysis in this section, it is not difficult to notice that always learning to classify adversarial samples to the same class as the corresponding benign ones is likely to be a too-hard assumption. In that case, trying to compulsively unify the learning of benign and adversarial samples, which may be two essentially inconsistent optimization objectives, becomes a logical and reasonable source of the existing trade-off between clean accuracy and adversarial robustness. This inspired us to propose a new AT paradigm in which the overstrict current assumption can be relaxed.
2.2 New Adversarial Training Paradigm with Dummy Classes
@@
A natural idea to relax the current assumption is to introduce additional dummy classes to accommodate the adversarial samples that are hard to classify to the original class, based on which we propose our new AT paradigm. In this section, we first provide preliminaries w.r.t. relevant concepts, followed by a formal formulation of our paradigm. Then via a toy case deliberately causing overfitting to the training adversary, we demonstrate the necessity to use soft labels than hard ones.

### v2

# New Paradigm of Adversarial Training: Releasing Accuracy-Robustness Trade-Off via Dummy Class

@@ Abstract
Adversarial Training (AT) is one of the most effective methods to enhance the robustness of Deep Neural Networks (DNNs). However, existing AT methods suffer from an inherent accuracy-robustness trade-off. Previous works have studied this issue under the current AT paradigm, but still face over 10% accuracy reduction without significant robustness improvement over simple baselines such as PGD-AT. This inherent trade-off raises a question: Whether the current AT paradigm, which assumes to learn corresponding benign and adversarial samples as the same class, inappropriately mixes clean and robust objectives that may be essentially inconsistent. In fact, our empirical results show that up to 40% of CIFAR-10 adversarial samples always fail to satisfy such an assumption across various AT methods and robust models, explicitly indicating the room for improvement of the current AT paradigm. To relax from this overstrict assumption and the tension between clean and robust learning, in this work, we propose a new AT paradigm by introducing an additional dummy class for each original class, aiming to accommodate hard adversarial samples with shifted distribution after perturbation. The robustness w.r.t. these adversarial samples can be achieved by runtime recovery from the predicted dummy classes to the corresponding original ones, without conflicting with the clean objective on accuracy of benign samples. Finally, based on our new paradigm, we propose a novel DUmmy Classes-based Adversarial Training (DUCAT) method that concurrently improves accuracy and robustness in a plug-and-play manner only relevant to logits, loss, and a proposed two-hot soft label-based supervised signal. Our method outperforms state-of-the-art (SOTA) benchmarks, effectively releasing the current trade-off. The code is available at https://github.com/FlaAI/DUCAT.
@@ 1 Introduction
Deep Neural Networks (DNNs) demonstrate remarkable performance in various real-world applications, but remain vulnerable to adversarial attacks (Biggio et al., 2013; Szegedy et al., 2014). Specifically, malicious input perturbations that are imperceptible to humans can cause significant changes in the output of DNNs (Goodfellow et al., 2015), raising serious security concerns within both the public and research communities (Wang et al., 2020b). In response, several defense mechanisms have been proposed to enhance the adversarial robustness of DNNs, such as defense distillation (Papernot et al., 2017), feature squeezing (Xu et al., 2018), randomization (Xie et al., 2018), and input denoising (Guo et al., 2018; Liao et al., 2018). However, most of them have proven subsequently to rely on obfuscated gradients (Athalye et al., 2018) and be ineffective against advanced adaptive attacks (Tramer et al., 2020), such as Auto-Attack (AA) (Croce and Hein, 2020b).
Currently, Adversarial Training (AT) (Goodfellow et al., 2015; Madry et al., 2018) is demonstrated as one of the most effective approaches to train inherently robust DNNs (Athalye et al., 2018; Dong et al., 2020). Different from clean training, the current AT paradigm directly learns adversarially augmented samples with the original labels to yield robust models. Research indicates that this paradigm is equivalent to optimizing an upper bound of natural risk on the original data, and can thereby serve as a principle against adversarial attacks (Tao et al., 2021). To the best of our knowledge, despite considerable advancements in the specific mechanism of existing AT methods, most of them remain adhering to this basic paradigm. The solid AT benchmarks such as PGD-AT (Madry et al., 2018), TRADES (Zhang et al., 2019) and MART (Wang et al., 2020b) are representative examples.
Figure 1: Conceptual difference between conventional AT and ours. The current AT paradigm assumes that the adversarial sample $\mathbf{x}^{\prime}$ should be assigned to the same class as the benign sample $\mathbf{x}$, which may be overstrict. In this work, we suggest introducing more dummy classes one-to-one corresponding to original ones, so that some hard $\mathbf{x}^{\prime}$ with shifted distribution can be accommodated without significantly hurting clean learning on $\mathbf{x}$.
However, recent studies indicate that existing AT methods suffer from an inherent trade-off between adversarial robustness and clean accuracy (Tsipras et al., 2019; Zhang et al., 2019; Raghunathan et al., 2019, 2020; Wang et al., 2020a; Bai et al., 2021). This means improving robustness via these AT methods is at the cost of reducing model accuracy compared with the standard training, which can seriously hurt the experience of benign users and greatly reduce the use of AT among real-world DNN application providers. A widely recognized reason of this problem is that AT equally requires predictive consistency within the $\epsilon$-ball of each sample, which can complicate decision boundaries, particularly for those samples near class margins, ultimately degrading clean generalization (Dong et al., 2022; Rade and Moosavi-Dezfooli, 2022; Cheng et al., 2022; Yang and Xu, 2022). To learn robustness w.r.t. such samples more appropriately, FAT (Zhang et al., 2020) and HAT (Rade and Moosavi-Dezfooli, 2022) utilized adaptive perturbations in AT to smooth decision boundaries; Cons-AT (Tack et al., 2022) encouraged the similarity of predictive distributions between adversarial samples derived from different augmentations of the same instance, thereby enhancing learnable patterns and reducing label noise; SOVR (Kanai et al., 2023) explicitly increased logits margin of important samples by switching from cross-entropy to a new one-vs-the-rest loss.
Despite several previous efforts, there has been limited substantial progress in addressing this trade-off problem. Most impressive advancements in recent years come from those methods introducing extra data (Alayrac et al., 2019; Carmon et al., 2019; Najafi et al., 2019; Li et al., 2022) or utilizing generative models (Wang et al., 2023b). However, they not only demand more resources and computational costs but also violate the conventional fairness assumption of AT (i.e., no additional data should be incorporated (Pang et al., 2021)). Other advanced AT techniques apart from these still typically experience a drop in clean accuracy exceeding 10% to date, and the state-of-the-art (SOTA) robustness has just improved slowly (i.e., about 1% per year on average since 2018) (Wei et al., 2023; Dong et al., 2023; Li and Spratling, 2023; Jia et al., 2024). This inherent trade-off raises a question about the current AT paradigm on uniformly learning accuracy and robustness:
Is the current AT paradigm, which compels DNNs to learn to classify the corresponding benign and adversarial samples into the same class, really appropriate and necessary for achieving robustness?
@@
In this work, as straightforward evidence supporting our deduction, we reveal that certain samples that always fail to meet the above objective of the current AT paradigm widely exist across various AT methods and different robust models. This observation implies that the conventional assumption of categorizing adversarial samples into the same class as the corresponding benign ones may be overstrict. As such, with the learning of benign and adversarial samples as two essentially inconsistent targets, the attempt to unify them in the current AT paradigm can be improper, which is likely to be blamed for the existing trade-off between clean accuracy and adversarial robustness. In response, we propose a new AT paradigm introducing additional dummy classes for certain adversarial samples that differ in distribution from the original ones to relax the current overstrict assumption. Then, we propose a novel AT method called DUmmy Classes-based Adversarial Training (DUCAT) accordingly, releasing the inherent trade-off between clean accuracy and robustness.
@@
Our core idea is to create dummy classes with the same number as the original ones and respectively attribute them as the primary targets for benign and adversarial samples during training. Importantly, we do not suggest a strict separation between corresponding benign and adversarial samples (i.e., to make the original classes completely benign and the dummy ones completely adversarial), because assuming corresponding benign and adversarial samples have completely different distributions is still excessive, which may cause memorization of hard-label training samples, resulting in overfitting to specific adversaries as shown by our toy case in Section 2.2.3. Instead, we construct unique two-hot soft labels to explicitly bridge corresponding original and dummy classes as the suboptimal alternative target of each other, so that the separation also becomes learnable, to utilize the potential of DNNs further. During inference time, such one-to-one correspondences enable the detection and recovery of adversarial samples, thereby achieving robustness without significantly compromising clean learning on benign samples. Specifically, if a test sample is classified into a dummy class, this serves as an indication of a potential adversarial attack, allowing us to recover its clean prediction through a projection back to the corresponding original class. It can then be viewed as the final prediction of the DUCAT-defended robust classifier. Separated from the computation graph of DNN and implemented in a run-time-only manner, such a projection further degrades the ability of real-world attackers.
Contributions. 1) For the first time, we empirically reveal that always-failed samples widely exist in conventional AT, explicitly suggesting the assumption of the current AT paradigm to learn benign and adversarial samples with the same labels is overstrict; 2) We propose a new AT paradigm introducing dummy classes to relax the current assumption, releasing the inherent accuracy-robustness trade-off caused by it, and correspondingly suggest a new two-hot soft label-based learning strategy for it; 3) A novel plug-and-play DUCAT method is proposed, concurrently improving the accuracy and robustness of four common AT benchmarks in large-scale experiments on CIFAR-10, CIFAR-100 and Tiny-ImageNet, significantly outperforming 18 existing SOTAs.
@@ 2 Involving Dummy Classes in Adversarial Training
2.1 Motivation: Always-Failed Cases Widely Exist in Adversarial Training
The undesirable progress regarding the trade-off between clean accuracy and adversarial robustness in the past years makes us suspect that the assumption of the current AT paradigm to assign corresponding benign and adversarial samples to the same class may be inappropriate and unnecessary. As intuitive evidence of this deduction, by two proof-of-concept experiments, we demonstrate that up to 40% adversarial samples that always fail to meet such an assumption generally exist in conventional AT crossing various AT methods and different robust models. The experiments are conducted on CIFAR-10 (Krizhevsky and Hinton, 2009) and ResNet-18 (He et al., 2016), with, as mentioned in Section 1, the most popular AT benchmarks, PGD-AT, TRADES, MART, and a representative SOTA, Consitency-AT. The training details are the same as our main experiments in Appendix C.1. For adversary, we adopt a typical and effective adversarial attack, PGD-10 (Madry et al., 2018).
Figure 3: High overlap between adversarial samples evading four AT benchmarks, implying such failures are more likely from an inappropriate learning objective under the current paradigm, rather than any specific AT methods.
@@ 2.1 Motivation: Always-Failed Cases Widely Exist in Adversarial Training
We first demonstrate that there is a high overlap in both successful and unsuccessful cases of the four AT methods. Specifically, for every benign test sample, with a high probability (i.e., $>$ 80%), its corresponding adversarial sample can either be uniformly defended by the four methods or escape from all of them. As shown in Figure 3, we report the numbers of samples that can respectively beat 0, 1, 2, 3, 4 (all) experimental defenses. The samples on which all the defenses are uniformly worked or failed make up the overwhelming majority. Besides, the confusion matrixes of the models protected by these four AT methods under the adversary are shown in Figure 9 in Appendix E, where similar failure patterns can be noted crossing all of them. Based on these observations, we state that always-failed adversarial samples widely exist no matter the specific choice of AT methods.
@@
By respectively adopting the four AT benchmarks to train a model from scratch and then test it by the same white-box adversary, above we have revealed that the always-failed cases are independent of specific AT methods. In this part, through black-box transfer attacks, we further demonstrate that such cases should also not be attributed to the weakness of any specific robust models. To be specific, we first train four different robust models respectively by the benchmarks and select any one of them as the surrogate model, based on which we generate adversarial samples for the transfer attack. Then we divide the generated samples into two subsets based on whether successfully attack the surrogate model, and respectively use them to attack the other three models. Figure 4 shows the results with the four robust models serving as the surrogate one in order, where significant differences in transferability between the originally successful and unsuccessful subsets can be observed. This means for every single adversarial sample generated based on any robust model, if it is originally effective/ineffective in attacking this surrogate model, then with a high probability, it would also work/fail on other robust models. This implies that always-failed cases generally exist regardless of the specific robust model as the target.
Figure 4: Robust models already enhanced by different AT methods are still highly likely to be uniformly beaten by the successful adversarial samples generated based on any one of these models. Such a deconstruction of adversarial transferability between robust models reveals the model-independent vulnerability of certain samples, which further supports our deduction for the current AT paradigm.
Figure 5: The logic flow of our motivation to introduce dummy classes in the new AT paradigm.
@@
Now that the widespread existence of always-failed samples is confirmed crossing different AT methods and robust models, compared with simply attributing this to specific technology or implementation details, it seems more reasonable to rethink whether it is the paradigm to be blamed in essence. Based on the analysis in this section, it is not difficult to notice that always learning to classify the adversarial samples to the same class as the corresponding benign ones is likely to be a too-hard assumption. In that case, trying to compulsively unify the learning of benign and adversarial samples, which may be two essentially inconsistent optimization objectives, becomes a logical and reasonable source of the existing trade-off between clean accuracy and adversarial robustness. The logic of our motivation is summed up with a flow chart as Figure 5, which inspired us to propose a new AT paradigm in which the overstrict current assumption can be relaxed via dummy classes.
2.2 New Paradigm of Adversarial Training with Dummy Classes
@@
A natural idea to relax the current assumption is to introduce additional dummy classes to accommodate the adversarial samples that are hard to classify into the original class, based on which we propose our new AT paradigm. In this section, we first provide preliminaries w.r.t. relevant concepts, followed by a formal formulation of our paradigm. Then, through a toy case deliberately causing overfitting to the training adversary, we demonstrate the necessity of designing two-hot soft labels rather than adopting conventional hard ones.